Solve the given equation:
2+6e^(4x)=19

Answer by nerdybill(7384)

2+6e^(4x)=19
Subtracting 2 from both sides:
6e^(4x)=17
Dividing both sides by 6:
e^(4x)=17/6
Take ln of both sides:
4x= ln(17/6)
Divide both sides by 4:
x= ln(17/6)/4
x = 0.2604